# appendices

### v1

@@ Methods
We use freely available SAR data from the Copernicus Sentinel-1 mission [19], consisting of two satellites equipped with C-band SAR instruments and moving in the same orbital plane, phased 180° apart. This setup originally allowed for global coverage with a revisit period below 6 days anywhere on Earth. Unfortunately this period has doubled to up to 12 days since the technical failure of Sentinel-1B in December 2021. For our study, we only use data from Sentinel-1A to maintain consistent temporal resolution. We utilize the Ground Range Detected (GRD) product available on GEE, consisting of log-amplitudes for the VV and VH polarizations, resampled at 10 m. Internally, GEE automatically preprocesses every tile with precise orbit correction, border and thermal noise removal, radiometric calibration, and terrain correction. We do not add any further preprocessing.
@@
For each location marked by UNOSAT, we stack the corresponding Sentinel-1 tiles and extract all backscatter values, for both the VV and VH polarizations. Limiting the training to the points annotated by UNOSAT corresponds to a fully supervised learning scheme where the target value is known for every training example. As the annotated points may be slightly offset due to large nadir angles or inaccurate annotations, we average the backscatter over a 3$\times$3 pixel window centered at the marked location. By relying only on this minimal neighborhood we disregard spatial context. We found that the time series signal carries most of the information relevant to our analysis, presumably because conflict-induced building damages are mostly small and localized. Per-pixel processing also scales particularly well on parallel processing architectures like the one underlying GEE.
@@
where $t_{\text{invasion}}$ is 2022-02-24, the day the invasion was launched; and $t_{\text{unosat}}$ is the acquisition date of the post-event image used for annotation by UNOSAT. All time series with $y_{i}=-1$ are discarded because we cannot determine whether they contain the damage event. We transform $x_{i}$ into $\phi(x_{i})$ using the following seven summary statistics: minimum, maximum, mean, median, standard deviation, kurtosis, and skewness. Finally, we configure the Random Forest to use 100 decision trees, with minLeafPopulation set to 3 and maxNodes set to 10,000. These values ensure sufficient robustness while remaining within the computational budget permitted by GEE.
@@ Appendix B Detailed Comparison with the PWTT Baseline
Label Precision Recall F1 F0.2 Accuracy AUC
Damaged 61.9 76.3 68.4 62.4 75.5 75.7
Undamaged 85.6 75.1 80 85.2
@@
Label Precision Recall F1 F0.2 Accuracy AUC
Damaged 69.7 78.7 73.9 70.0 80.7 80.3
Undamaged 87.8 81.8 84.7 87.6
Table 6: Quantitative performance of our approach on our test set, with confidence threshold 0.5. All metrics are given as percentages.
Appendix C Choice of optimal threshold
There is an inherent trade-off between precision and recall, and the optimal threshold depends on the specific application. To minimize false alarms in our analysis, we optimized the F0.2-score for the figures presented in this paper, and found an optimal confidence threshold of 0.64. In practice, users of our tool can dynamically adjust the confidence threshold to suit their specific needs. Fig. 6 illustrates how our metrics vary with different thresholds.
Figure 6: Pixel-level performance of our method on our test set, for different confidence thresholds. We find that a confidence threshold of 0.64 optimizes the F0.2-score.
@@ Appendix D Temporal Evolution of Damage
Figure 7: Percentage of buildings likely damaged for each period, cumulated since the beginning of the war and aggregated by hromadas. The predictions were thresholded at 0.64, and only buildings larger than 50 m2 were considered.
@@ Appendix E Limitations of Building Footprints
The building footprints provided by Overture Maps, which for Ukraine were compiled from OpenStreetMap and Microsoft, inevitably come with limitations. For instance, OpenStreetMap data may be obsolete due to the demolition of older structures (Fig. 8.A) or the construction of new ones (Fig. 8.B). While Microsoft footprints serve as a valuable complement to OpenStreetMap data, they were automatically retrieved with a deep neural network from Bing Maps imagery, which may not always offer recent images, and suffers from gaps due to cloudy or outdated acquisitions as for example in Kherson (Fig. 8.C). Moreover, it is not uncommon for large non-building structures, e.g. containers (Fig. 8.D), boats (Fig. 8.E) or even aircraft (Fig. 8.F), to be wrongly identified as buildings.
@@ Appendix F World Bank Estimates
The World Bank’s third Rapid Damage and Needs Assessment estimated that 547,000 buildings have been completely destroyed in the Ukrainian conflict as of December 31, 2023 [2, p. 76], which aligns with our estimates of over 500,000 buildings damaged over the first two years of conflict.
@@
Figure 9: Comparison of the World Bank damage distribution by oblasts against our estimates.

### v2

@@ Methods
We use freely available SAR data from the Copernicus Sentinel-1 mission [39], consisting of two satellites equipped with C-band SAR instruments and moving in the same orbital plane, phased 180° apart. This setup originally allowed for global coverage with a revisit period below 6 days anywhere on Earth. Unfortunately, this period has doubled to up to 12 days since the technical failure of Sentinel-1B in December 2021. For our study, we only use data from Sentinel-1A to maintain consistent temporal resolution. We utilize the Ground Range Detected (GRD) product available on GEE, consisting of log-amplitudes for the VV and VH polarizations, resampled at 10 m. Internally, GEE automatically preprocesses every tile with precise orbit correction, border and thermal noise removal, radiometric calibration, and terrain correction. We do not add any further preprocessing.
@@
For each location marked by UNOSAT, we stack the corresponding Sentinel-1 tiles and extract all backscatter values, for both the VV and VH polarizations. Limiting the training to the points annotated by UNOSAT corresponds to a fully supervised learning scheme where the target value is known for every training example. Relying only on pixel-wise time series discards all spatial context. We found that the time series signal carries most of the information relevant to our analysis, presumably because conflict-induced building damages are mostly small and localized. Per-pixel processing also scales particularly well on parallel processing architectures like the one underlying GEE.
@@
where $t_{\text{invasion}}$ is 2022-02-24, the day the invasion was launched; and $t_{\text{unosat}}$ is the acquisition date of the post-event image used for annotation by UNOSAT. All time series with $y_{i}=-1$ are discarded because we cannot determine whether they contain the damage event. We transform $x_{i}$ into $\phi(x_{i})$ using the following seven summary statistics: minimum, maximum, mean, median, standard deviation, kurtosis, and skewness. Finally, we configure the Random Forest to use 50 decision trees, with minLeafPopulation set to 3 and maxNodes set to 10,000. These values ensure sufficient robustness while remaining within the computational budget permitted by GEE.
@@
Code Availability
All the code necessary to reproduce our results can be found in the following repository: https://github.com/olidietrich/ukraine-damage-mapping-tool/.
Data Availability
All data used in this study is publicly accessible. Links to all results, including damage heatmaps and building footprints with corresponding damage estimates, will be made available in the repository hosting our code.
@@ Appendix B Detailed Comparison with the PWTT Baseline
Label Precision Recall F1 Accuracy AUC
Damaged 61.9 76.3 68.4 75.5 75.7
Undamaged 85.6 75.1 80
@@
Label Precision Recall F1 Accuracy AUC
Damaged 67.1 84.6 74.9 80.3 81.3
Undamaged 90.5 78.0 83.8
Table 6: Quantitative performance of our approach on our test set, with confidence threshold 0.5 (reprinted from Table 1). All metrics are given as percentages.
Appendix C Choice of threshold
There is an inherent trade-off between precision and recall, and the optimal threshold depends on the specific application. To keep false alarms at a low enough rate, we set a target precision of 90%, for which we find an optimal confidence threshold of 0.655. That value has been used to generate all metrics in the present paper unless stated otherwise. With this threshold, the user can expect a recall of 46.5%. In practice, users of our tool can dynamically adjust the confidence threshold to suit their specific needs. For instance, one could start with a high threshold to discover hotpots with a very high likelihood of destruction, then gradually decrease the threshold to increase the recall and better assess the extent of destruction. Fig. 6 illustrates how our metrics vary with different thresholds.
Figure 6: Pixel-level performance of our method on our test set, for different confidence thresholds. We set a target desired precision of 90% and found a confidence threshold of 0.655. The two vertical lines indicate the thresholds 0.5 and 0.655.
@@ Appendix D Temporal Evolution of Damage
Figure 7: Percentage of buildings likely damaged for each period, cumulated since the beginning of the war and aggregated by hromadas. The predictions were thresholded at 0.655, and only buildings larger than 50 m2 were considered.
@@ Appendix E Limitations of Building Footprints
The building footprints provided by Overture Maps, which for Ukraine were compiled from OpenStreetMap and Microsoft, inevitably come with limitations. For instance, OpenStreetMap data may be obsolete due to the demolition of older structures (Fig. 8.A) or the construction of new ones (Fig. 8.B). While Microsoft footprints serve as a valuable complement to OpenStreetMap data, they were automatically retrieved with a deep neural network from Bing Maps imagery, which may not always offer recent images, and suffers from gaps due to cloudy or outdated acquisitions, for example in Kherson (Fig. 8.C). Moreover, it is not uncommon for large non-building structures, e.g. containers (Fig. 8.D), boats (Fig. 8.E) or even aircraft (Fig. 8.F), to be wrongly identified as buildings.
@@ Appendix F World Bank Estimates
In this section, we compare our estimates, with a total of $\approx$400,000 buildings damaged between Feb 2022 to Feb 2024, to figures provided by the World Bank (WB) in their third Rapid Damage and Needs Assessment report [2, p. 74-80].
According to the WB report, a total of 2,106,920 housing units were damaged by December 2023. Among these, 547,010 were destroyed, 679,382 sustained medium damage and 880,528 suffered minor damage. Importantly, the WB figures refer to housing units (HU), not buildings. A single-family home counts as one HU, but multi-family buildings may contain dozens of HU, making a direct comparison to our building-based estimates difficult.
@@
Fig. 9 shows the differences between our estimates and those of the WB. Excluding the other oblasts outside the direct conflict zone, both distributions align quite well, with indeed the majority of damage concentrated in Donetska and Kharkivska. However, our model estimates significantly more damage than the WB for the other oblasts, i.e. the ones further from the frontline. It appears that construction and other non-conflict-related changes dominate in those regions, which our model is unable to distinguish from building damages. This highlights that there is a trade-off between, on the one hand, comprehensive screening to detect also unexpected conflict-induced destruction far from the known frontline; and, on the other hand, accurate assessment by ruling out areas that are a-priori unlikely to be affected.
@@
Figure 9: Comparison of the geographical distribution of the World Bank estimates against ours. The World Bank provides spatial distribution for their damage cost estimates, while our spatial distribution shows buildings likely damaged.
Appendix G Changes in Backscatter due to Destruction
The backscatter signal from a building is influenced by numerous variables, including building characteristics (e.g., geometry, orientation, surface texture, material composition), satellite parameters (e.g., frequency, polarization, incidence angle, orbit direction), and environmental factors (e.g., vegetation, soil moisture, snow, or heavy rainfall). Intuitively, one might expect that intact structures yield stronger specular returns, with flat surfaces and right-angle corners acting as corner reflectors, whereas structural damage would disrupt these patterns, producing a more diffuse signal. However, several studies indicate that damage signatures in SAR data are highly variable and depend on the type and extent of destruction [19, 20]. Damaged buildings can even exhibit backscatter signals as high as undamaged structures, or even higher [53].
Our experiments corroborate these findings. Fig. 10 presents pixel-wise backscatter time series that illustrate how conflict-related destruction appears in the data. High-resolution pre- and post-destruction satellite imagery from Google Earth is provided for reference. We observe that the backscatter amplitude following destruction can vary widely – increasing, decreasing, or remaining largely stable – depending on the location, the satellite orbit, and the polarization. These examples emphasize the need to utilize all available orbits in order to maximize reliability.
Figure 10: Sentinel-1 backscatter time series at selected UNOSAT locations for VV and VH polarizations, with VHR satellite images from Google Earth showing pre- and post-destruction conditions. Vertical dashed lines indicate the date of Russia’s full-scale invasion (black) and the dates of the displayed pre-destruction (green) and post-destruction (red) images. For readability, only two orbits per location are shown, even if in some locations up to four orbits are available. Amplitude can decrease (B, C), increase (A, C) or remain relatively stable (B, D) after destruction. The same event may manifest differently depending on the flight direction (B), or even on the incidence angle within the same flight direction (C).
Appendix H Ablation studies
We have performed an ablation study to support our design choices and understand the impact of various settings on our model. We kept the same train/test split and label definitions as described in Evaluation. To reduce computational cost, we opted for a pixel-wise comparison to the UNOSAT labels, contrary to Evaluation, where we ran the inference densly over the entire area of interest and allowed for shifts up to $\pm$1 pixel in $x$ and $y$ to account for geo-location uncertainty of the labels.
The $F_{1}$-score of our optimal configuration, which utilizes both VV and VH Sentinel-1 bands, all seven features, 50 decision trees and extracts the time series pixel-wise, is 0.72. Fig. 11 summarized the outcome for different settings. We observe the following characteristics:
Input bands: Our tests confirmed that VV polarization provides more relevant information for detecting destruction compared to VH. However, using both bands together produced the best results, indicating that they complement each other to some degree.
Input data: We tried replacing Sentinel-1 SAR data with Sentinel-2 optical images [54]. Pixel-wise Sentinel-2 time series were extracted from the Level-2A collection available on GEE [55], from which we filter out cloudy pixels using GEE’s CloudScore+ mask [56]. Various band combinations were tried, but incorporating Sentinel-2 data never enhanced model performance. This was expected, as our pipeline was design for Sentinel-1 properties, such as stable pre-event signals and a focus on temporal context over spatial context. Intuitively, it is harder to distinguish intact buildings from damaged ones by looking at the brightness change at an individual pixel location. Also combining Sentinel-1 and Sentinel-2 led to poorer results than using Sentinel-1 alone, likely due to the confounding effect of a six-fold increase in features, most of which contain little information for the task.
Model: We evaluated alternative classifiers available in GEE, i.e., Boosted Trees [52] and SVM [57]. Boosted trees showed weaker performance, while taking approximately 3$\times$ longer to run. We were unable to obtain results for the SVM: when applied to our data it quickly exceeded the computational budget allowed by GEE.
Number of trees: We tested models with 10, 25, 50, 75, and 100 trees. Performance gains were minimal beyond 50 trees, so we chose 50 as the optimal number, balancing accuracy and computational cost. Above 100 trees, we again hit GEE’s bound on the compute budget.
Features: We tried different configurations, from a simple combination of the mean and standard deviation to the entire set of seven features. We observed consistent gain as the number of features increased.
Extraction window: Instead of extracting the time series pixel-wise, we experimented with taking the spatial mean over a 3$\times$3 neighborhood to suppress noise. The resulting heatmaps are smoother, but this comes at the cost of missing many small-scale and sparse damage events.
Figure 11: Ablation studies. $F_{1}$-score for different configurations.